OpenStudy (anonymous):

The speed of a passenger train is 12mph faster than the speed of a freight train. The passenger train travels290 miles in the same time it takes the freight train to travel 230 miles. Find the speed of each train. Freight train=? Passenger train=?

OpenStudy (anonymous):

Let the speed of frt train = x mph then speed of passng train = (x+12) mph let the equal time for which they travel be = y hours Distance coverd by frt train in y hours = speed x time = xy we are given xy = 230 ----(A) distance covered by passng train = (x+12)*y = xy + 12y we are given xy + 12y = 290 -----(B) substituting xy=230 from (A) in (B), we get 230 +12y = 290 12y = 290-230 y=60/12 = 5 Putting y=5 in (A) we get 5x= 230 x = 230/5 = 46 mph so speed of frt train is 46 mph and speed of passngr train = 46+12 = 58 mph

OpenStudy (anonymous):

if u r happy with the answer, pls click GOOD ANSWER button for me ( :})

OpenStudy (anonymous):

Let the speed of frt train = x mph then speed of passng train = (x+12) mph let the equal time for which they travel be = y hours Distance coverd by frt train in y hours = speed x time = xy we are given xy = 230 ----(A) distance covered by passng train = (x+12)*y = xy + 12y we are given xy + 12y = 290 -----(B) substituting xy=230 from (A) in (B), we get 230 +12y = 290 12y = 290-230 y=60/12 = 5 Putting y=5 in (A) we get 5x= 230 x = 230/5 = 46 mph Speed of frt train is 46 mph Speed of passnger train = 46+12 = 58 mph

OpenStudy (anonymous):

copying and pasting not appreciated ( :))